### v3

# Assessing the Reasoning Capabilities of LLMs in the context of Evidence-based Claim Verification

Abstract
Although LLMs have shown great performance on Mathematics and Coding related reasoning tasks, the reasoning capabilities of LLMs regarding other forms of reasoning are still an open problem. Here, we examine the issue of reasoning from the perspective of claim verification. We propose a framework designed to break down any claim paired with evidence into atomic reasoning types that are necessary for verification. We use this framework to create RECV, the first claim verification benchmark, incorporating real-world claims, to assess the deductive and abductive reasoning capabilities of LLMs. The benchmark comprises of three datasets, covering reasoning problems of increasing complexity. We evaluate three state-of-the-art proprietary LLMs under multiple prompt settings. Our results show that while LLMs can address deductive reasoning problems, they consistently fail in cases of abductive reasoning.Moreover, we observe that enhancing LLMs with rationale generation is not always beneficial. Nonetheless, we find that generated rationales are semantically similar to those provided by humans, especially in deductive reasoning cases.
Assessing the Reasoning Capabilities of LLMs in the context of Evidence-based Claim Verification
John Dougrez-Lewis1,∗  Mahmud Elahi Akhter2,∗  Federico Ruggeri3 Sebastian Löbbers2Yulan He4,5Maria Liakata2,5 University of Warwick, UK1 Queen Mary University of London, UK2 University of Bologna, Italy3 King’s College London, UK4 The Alan Turing Institute, UK5 j.dougrez-lewis@warwick.ac.uk, yulan.he@kcl.ac.uk, federico.ruggeri6@unibo.it {m.akhter, s.lobbers, m.liakata}@qmul.ac.uk These authors contributed equally to this work.
1 Introduction
Large Language Models (LLMs) have shown remarkable proficiency in complex tasks where reasoning capabilities, such as logical deduction and semantic comparison, are paramount. Notable examples include solving MBA exams Terwiesch (2023), passing professional medical tests Kung et al. (2023); Nori et al. (2023), performing quantitative reasoning Lewkowycz et al. (2022), and communication games Bakhtin et al. (2022); Xu et al. (2023); Gandhi et al. (2023). However, there is ongoing debate about whether such proficiency is due to LLMs manifesting reasoning capabilities or rather pattern matching and semantic similarity via memorization. For example, earlier claims that LLMs posses Theory of Mind (ToM) capabilities Bubeck et al. (2023); Kosinski (2023) were shown to be inaccurate Ullman (2023); Sileo and Lernould (2023). In particular, despite appearing to manifest some form of ToM capabilities, LLMs mostly rely on shallow heuristics and spurious correlations Shapira et al. (2023). Additionally, preliminary observations of emergent reasoning capabilities Wei et al. (2022) were subsequently attributed to metric choice Schaeffer et al. (2023), in-context learning Lu et al. (2023b), and shortcuts Kavumba et al. (2019).
These findings motivate the need for further research on the reasoning capabilities of LLMs, especially in high-stake real-world applications, where research on this topic is in its infancy. A notable example is fact-checking, where LLMs are considered to hold great potential for increased productivity overshadowed by the ease with which bad actors can proliferate misinformation Guo et al. (2023). Verifying information is challenging since models require both accurate classification and strong rationale generation to be effective Schlichtkrull et al. (2023). It is thus essential to understand the reasoning capabilities and limitations of LLMs in the context of fact-checking. In particular, we extend the current discussion around the reasoning abilities of LLMs, focusing on their ability to verify real-world claims.
In this work, we first propose a framework for breaking down complex claims into atomic reasoning steps. The motivation behind this is the lack of uniform terminology around reasoning evaluation. Most prominent evaluation datasets for reasoning are based on mathematics and coding, which involves deductive reasoning but is not treated as such  Sprague et al. (2024). Our framework is rooted in existing philosophy literature concerning logical reasoning that aligns well with NLP Wason and Johnson-Laird (1972); Galotti (1989).
We use our framework to create Reasoning in Evidence-based Claim Verification (RECV), the first reasoning benchmark for claim-verification. The benchmark comprises three datasets, curated from existing resources targeting different domains: VitaminC Schuster et al. (2021) from Wikipedia, CLIMATE-FEVER Diggelmann et al. (2020) from online claims and Wikipedia, and PHEMEPlus Dougrez-Lewis et al. (2022) from rumours circulating on social media and associated evidence from news articles. The claims involve increasing levels of complexity as we move from VitaminC to PHEMEPlus, often requiring deductive and/or abductive reasoning.
We use RECV to evaluate three state-of-the-art proprietary LLMs that have shown impressive performance on various reasoning and language benchmarks Huang and Chang (2023); DeepSeek-AI et al. (2025). These models are: Claude V3 Sonnet Anthropic (2023), GPT-4 OpenAI (2023), and GPT-4o OpenAI et al. (2024). In particular, we prompt models with and without Chain-of-Thought (CoT) Wei et al. (2023) rationale generation to assess if and how the latter influences reasoning. In alignment with previous work Saparov et al. (2023); Akyürek et al. (2024); Li et al. (2024) we find that LLMs are capable of deductive reasoning. However, they consistently fail at claim verification when presented with evidence that requires abductive reasoning. Furthermore, we observe conflicting results when prompting LLMs with CoT strategies. In particular, CoT leads to performance improvements for simple claim verification as in VitaminC, but detrimental in the case of complex claims such as those found in CLIMATE-FEVER and PHEMEPlus. Lastly, we carry out a qualitative analysis of generated rationales and observe high semantic similarity with human explanations, especially in deductive reasoning cases.
In summary, we make the following contributions:
  • $\bullet$
    We propose a framework for decomposing claim-evidence pairs into atomic reasoning types for verification, covering deductive and abductive reasoning (§3).
  • $\bullet$
    We create the first reasoning benchmark for claim verification comprising three datasets of increasing complexity (§4).
  • $\bullet$
    We extensively evaluate the reasoning capabilities of three state-of-the-art LLMs, showing that models fail when it comes to abductive reasoning and CoT’s effectiveness is task-dependent (§5).
  • $\bullet$
    We show that generated rationales are consistent with human reasoning for correct predictions, but the model is often unable to leverage such rationales for claim verification (§6).
Figure 1: Resolution of claim verification via a single-step abductive reasoning type using RECV framework.
2 RECV Logical Framework
Reasoning is often used interchangeably to denote critical thinking, decision-making, and logical reasoning. Following Wason and Johnson-Laird (1972) and Galotti (1989), we define reasoning as the process of logical steps that result in some form of decision-making or conclusion. Thus we define reasoning as a series of inference steps linking claims and evidence to reach a conclusion.
In particular we consider that reasoning consists of the interplay of three interrelated components: types, processes, and tasks. This is the basis of our RECV framework. Reasoning types are different forms of logical inference that we can use to reach a conclusion from a set of observations or premises. We distinguish between atomic and compound reasoning types. Atomic types denote basic forms of logical inference and include deductive, abductive, inductive, and analogical reasoning.
A reasoning task is any task that requires multiple reasoning types, often in complex interaction with each other. For example claim verification is a composite reasoning task. A reasoning process is the method of interaction between reasoning types or even tasks within complex reasoning tasks. Notable examples of reasoning processes are multi-hop or multi-step inference, where individual steps or hops can be of different reasoning types. In this paper we focus particularly on atomic reasoning types.
2.1 Atomic Reasoning Types
Deduction:
A conclusion is drawn directly from evidence. In the context of a claim, if the evidence supports the claim then the claim is deduced to be true (if $P$ then $Q$, where $P$ is the claim and $Q$ is the evidence). For example,
  • $\bullet$
    $P$: Schools closed, Dammartin-en-Goele residents told to stay indoors, town ‘like warzone’. [Claim]
  • $\bullet$
    $Q$: Schools went into lockdown and the town appealed to residents to stay inside residents’ houses. [Evidence]
  • $\bullet$
    $C$: Here, $P\implies Q$. The schools have been closed and citizens have been told to stay home. Thus, the town is like in a warzone situation. [Conclusion]
Equally if the evidence contradicts the claim then the claim is deduced to be false.
  • $\bullet$
    $P$: Heart goes out to 148 passengers and crew of Germanwings Airbus A320 that has crashed in French Alps, Southern France. [Claim]
  • $\bullet$
    $Q$: German jetliner carrying 144 passengers and six crew en route from Barcelona, Spain, to Düsseldorf, Germany, has crashed in the French Alps, killing all 150 people on board. [Evidence]
  • $\bullet$
    $C$: Here, $Q$ contradicts $P$. The evidence directly states the death toll is 150 which refutes the claim. [Conclusion]
Abduction:
The most plausible conclusion is drawn from a set of candidate hypotheses, based on partial evidence. Abduction could lead to false conclusions.
  • $\bullet$
    Claim: Pluto’s climate change over the last 14 years is likely a seasonal event.
  • $\bullet$
    Evidence: The long orbital period of Neptune results in seasons lasting forty years. As a result, Neptune experiences similar seasonal changes to Earth. There’s evidence for methane escape and strong seasonal and dynamical perturbations of Neptune’s atmospheric temperatures. Each planet therefore has seasons, changes to the climate over the course of its year.
  • $\bullet$
    Conclusion: The evidence only mentions Neptune. However, the claim is regarding Pluto. Given the partial evidence, the claim is supported based on the plausible hypothesis that Pluto is near Neptune and it is likely to have similar attributes when it comes to seasons and climate change.
Induction:
An inference is drawn from complete evidence (in a specific domain) and then a generalization (a rule that can be used beyond the initial domain) is derived from it. As per Flach and Kakas (2000), for inductive reasoning, the evidence can be true whilst only providing partial support for the conclusion, which typically generalizes beyond the evidence itself. Such generalization indicates there is no guarantee that the conclusion is true elsewhere.
Analogical reasoning:
Conclusions are drawn based on the similarities between entities. While we do not provide examples for inductive and analogical reasoning in this section, they are still part of our framework. The focus on deduction and abduction is justified in (§3). We provide more formal definitions of atomic reasoning types in Appendix A with additional examples.
3 Methodology
We discuss our methodology for reasoning in claim verification. We first showcase the application of our RECV framework and then motivate our focus on deduction and abduction via a preliminary study.
RECV Logical Framework Application:
The application of the RECV framework can be seen in Figure 1. The reasoning task here is claim verification and the reasoning type is composite. The claim is resolved using a single-step process, that consists of abductive type atomic reasoning. Here we only highlight the most plausible hypothesis that resolves the claim as true. However, in practice, we would generate multiple hypotheses before coming to the most plausible one.
Preliminary Study:
Our objective here was to determine the atomic reasoning types necessary for accomplishing claim verification from evidence. We first collected a small dataset by manually selecting 90 claims and associated evidence from VitaminC Schuster et al. (2021), CLIMATE-FEVER Diggelmann et al. (2020), and PHEMEPlus Dougrez-Lewis et al. (2022). We focus on these resources as they are widely used in claim and rumour verification and differ in complexity.
Two annotators with expertise in Computer Science and native English proficiency assigned reasoning type labels to claim-evidence pairs following (§2). Disagreements encountered were resolved via a discussion stage with an independent expert. The Inter-Annotator Agreement (IAA) measured as Bennett’s S score Bennet et al. (1954) to account for label imbalance of reasoning types is 0.90, denoting almost perfect agreement. We observe that all examples either require deductive or abductive reasoning types.
Deductive and Abductive Reasoning:
Our preliminary investigation suggests that inductive and analogical reasoning are rarely employed in claim verification. This is presumably because inductive reasoning relies on complete evidence, which is rarely available in real-world domain-specific settings. Generalisations from one domain to another, relevant to inductive reasoning, may only occur in scenarios that share common background knowledge, as in the medical domain.
Similarly, analogical reasoning may be more suitable for other fact-checking related tasks like profiling and motive analysis where comparing information bits frequently occurs to reach a conclusion. By contrast, deductive and abductive reasoning types are often required in fact-checking Pan et al. (2023); Tan et al. (2024). For these reasons, here we focus on deduction and abduction. We show that they represent a challenging setting for claim verification (§4), and model evaluation with LLMs (§5).
4 RECV Benchmark
We discuss the creation of RECV, in particular, our sample selection strategy and data annotation. See Appendix B for details regarding the three datasets.
Data Sampling Strategy.
We build a heuristic-based sampling strategy to mitigate the anticipated data imbalance between deductive and abductive samples, as it was important to ensure both are represented in the annotated data.
We used a combination of three embedding-based text similarity metrics to compute the average claim similarity between deductive samples. Likewise for abductive samples. The metrics are: cosine similarity, BERTScore Zhang et al. (2020) and BLEURT score Pu et al. (2021). We used the data collected during our preliminary study (§3) to set a similarity threshold for each reasoning type. We used each threshold to sample claims likely to be resolved via deductive and abductive reasoning, respectively. In particular, we exclude instances labeled as ‘unverified’ since such claims are always associated with deductive reasoning, either due to lack of proper evidence or contradictory evidence. In total, we sampled 500 claim-evidence pairs from each dataset. See Appendix C for more details on data sampling.
VitaminC Supported Refuted Total Deductive 272 199 471 Abductive 11 18 29 Total 283 217 500 CLIMATE-FEVER Supported Refuted Total Deductive 269 129 398 Abductive 88 14 102 Total 357 143 500 PHEMEPlus Supported Refuted Total Deductive 336 128 464 Abductive 22 14 36 Total 358 142 500
Table 1: RECV statistics.
Data Annotation
We recruited 9 PhD students in Computer Science, fluent in English and grouped them in triples, one for each dataset. We evenly distributed dataset samples to annotators in a triple, so that 100 samples were annotated by all. In total, each annotator in a triple labeled 233 samples. Annotation guidelines per dataset are in Appendix D. We computed IAA as Bennett’s S score to account for label imbalance (see Appendix E for pairwise agreement scores). The IAA is 0.75 for VitaminC, 0.56 for CLIMATE-FEVER, and 0.67 for PHEMEPlus. Table 1 reports our RECV statistics.
In particular, we observe that the rate of abductive reasoning samples is relatively low compared to deductive ones: 5.8% in VitaminC, 20.4% in CLIMATE-FEVER, and 7.2% in PHEMEPlus. This imbalance is expected given the nature of collected evidence; most evidence provided, either in the form of Wikipedia articles as in VitaminC and CLIMATE-FEVER or news articles as in PHEMEPlus, contains detailed information to deductively verify the claim. In total, RECV consists of 1500 claim-evidence pairs with associated veracity and reasoning labels. The average sentence length for evidence in VitaminC is 1.084, 7.562 for PHEMEPlus and 7.828 for CLIMATE-FEVER. This highlights the varying complexity of the datasets and RECV.
5 Claim Verification with LLMs
Our objective here is to assess the capabilities of LLMs in performing deductive and abductive reasoning to determine the veracity of a claim.
Setup:
We formulate claim verification as a prediction task. Given a claim-evidence pair, we prompt LLMs to predict whether the evidence supports or refutes the claim (Figure A2 (bottom)). We consider two settings: No-Exp and Exp. In No-Exp, we prompt LLMs to predict the claim veracity without any rationale generation. In Exp, we first prompt LLMs to produce a rationale and then use the generated information to predict claim veracity. For each setting, we consider two different prompt strategies: Zero-Shot (ZS), and Manual Chain-of-Thought (M-CoT) Wei et al. (2023). In addition, in Exp, we also consider Zero-Shot Chain-of-Thought (ZS CoT) Kojima et al. (2023). ZS CoT was applied only under Exp as rationale generation is integral to ZS CoT prompting. In all the prompts, we provide dataset specific personas and instructions in the system prompt and CoT examples in the user prompt. We report the prompts in Appendix F.
Metrics:
We compute macro F1 score for veracity of claims given the evidence and the error rate of claim-evidence pairs concerning deductive and abductive reasoning types, respectively. F1 was chosen due to the class imbalance in CLIMATE-FEVER and PHEMEPlus (they have a 70/30 ratio between support and refute labels). We use annotators’ reasoning type labels for claim-evidence pairs to identify errors in verification per category (cases of abduction vs deduction) and express it via error rate.
Models:
We consider three state-of-the-art proprietary LLMs with remarkable proficiency in a wide range of tasks: Claude V3 Sonnet Anthropic (2023), GPT-4 OpenAI (2023), and GPT-4o OpenAI et al. (2024). We conducted our experiments using OpenAI and Anthropic’s official API.
5.1 Results
Table 2 reports classification performance and error rates per reasoning type for claim verification on RECV. We discuss dataset-specific results in detail.
VitaminC CLIMATE-FEVER PHEMEPlus Model F1 $\uparrow$ Deductive $\downarrow$ Abductive $\downarrow$ F1 $\uparrow$ Deductive $\downarrow$ Abductive $\downarrow$ F1 $\uparrow$ Deductive $\downarrow$ Abductive $\downarrow$ Claude ZS No-Exp 0.85 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.80 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.73 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ Claude M-CoT No-Exp 0.87 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.80 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.76 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ GPT-4 ZS No-Exp 0.86 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.87 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.69 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ GPT-4 M-CoT No-Exp 0.90 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.85 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.70 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ GPT-4o ZS No-Exp 0.88 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.84 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.72 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ GPT-4o M-CoT No-Exp 0.88 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.92 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.74 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ Claude ZS Exp 0.89 9.79(+3.83) 30.00(+3.33) 0.74 17.34(-4.53) 52.94(-12.75) 0.74 20.04(-0.65) 41.67(-2.78) Claude ZS CoT Exp 0.88 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.70 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.73 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ Claude M-CoT Exp 0.90 8.72(+4.04) 30.00(-6.67) 0.73 17.59(-4.77) 57.84(-16.01) 0.73 23.49(-4.96) 41.67(-2.78) GPT-4 ZS Exp 0.88 10.64(+1.49) 36.67(-6.67) 0.78 15.08(-6.28) 47.06(-26.47) 0.73 20.04(+0.65) 52.78(-13.89) GPT-4 ZS CoT Exp 0.88 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.77 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.71 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ GPT-4 M-CoT Exp 0.89 8.72(-0.42) 36.67(-10.00) 0.82 11.31(-1.26) 35.29(-7.84) 0.73 19.61(+2.80) 50.00(+2.78) GPT-4o ZS Exp 0.89 9.15(+1.28) 30.00(+10.00) 0.79 14.07(-4.52) 42.16(-8.82) 0.74 18.32(+1.72) 44.44(-2.78) GPT-4o ZS CoT Exp 0.89 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.78 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ 0.74 ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ ${}^{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ GPT-4o M-CoT Exp 0.89 8.94(+9.96) 36.67(-6.67) 0.78 13.82(-4.77) 42.16(-16.67) 0.75 18.75(+0.65) 47.22(+0.00)
Table 2: Claim verification performance on RECV. Best results are in bold, second-best results are underlined. We report error rate delta performance between No-Exp and Exp settings in brackets. Negative delta indicates that rationale generation degrades perforamnce.
VitaminC:
Among prompting strategies, M-CoT leads to the highest increase in performance across all models. The average error rate across all models and settings is $10.31\%$ for deductive reasoning and $32\%$ for abductive reasoning. This shows that all models struggle with abductive reasoning, even in less challenging settings like VitaminC. Regarding model settings, we observe conflicting results. In particular, generating rationales improves veracity classification performance for deductive samples in all models, except for GPT-4 M-CoT. By contrast, only Claude ZS and GPT-4o ZS show improvements in Exp compared to No-Exp when targeting abductive reasoning. Overall, when moving to the Exp settings, we observe a $7.5\%$ average performance drop, with GPT-4 reporting the highest degradation ($-10\%$). Lastly, regarding prompting strategies, we observe that M-CoT outperforms CoT in deductive cases, while reporting comparable results in abductive ones.
CLIMATE-FEVER:
The average error rate across all models and settings is $15.58\%$ for deductive reasoning and $48.58\%$ for abductive reasoning. Similar to VitaminC, these results denote that LLMs fail at predicting claim veracity when dealing with abductive reasoning. In particular, abductive reasoning samples are on average three times more challenging than deductive ones. Regarding model settings, we observe that rationale generation leads to performance degradation in all scenarios. Overall, we observe a $4.36\%$ average performance drop for deductive cases and $14.76\%$ for abductive ones. Regarding prompting strategies, we observe similar results to VitaminC where M-CoT outperforms CoT. In particular, the average error rate for M-CoT is $14.24\%$ on deductive cases ($+2.76$) and $45.1\%$ on abductive ones ($+8.17$).
PHEMEPlus:
The results suggest that there is no model or prompting strategy that consistently outperforms others. The average error rate across all models and settings is $20.06\%$ for deductive reasoning and $44.68\%$ for abductive reasoning. Compared to VitaminC and CLIMATE-FEVER, PHEMEPlus represents a more challenging setting for deductive reasoning, while it is comparable in complexity with CLIMATE-FEVER when assessing LLMs for claim verification. Regarding model settings, we observe minor performance improvements when prompting LLMs to generate rationales in deductive cases, with a $1.46\%$ average gain. Claude is the only exception with a $2.81\%$ average performance drop when moving to Exp. By contrast, we observe notable performance degradation in abductive reasoning cases, with GPT-4 ZS Exp being the worst ($-13.89\%$). Lastly, regarding prompting strategies, we observe no performance difference between ZS CoT and M-CoT, highlighting the higher task complexity in PHEMEPlus.
6 Explanation evaluation
Providing reasonable explanations to support predicted veracity labels is a crucial aspect of claim verification systems. In particular, an automated system needs to be both convincing and trustworthy to convince users in practice Schlichtkrull et al. (2023). Therefore, we evaluate the LLMs generated rationales in the Exp setting. This is paramount considering that LLMs tend to hallucinate Bouyamourn (2023); Rawte et al. (2023) and be self-contradictory at times Mündler et al. (2023). We randomly selected 100 samples from each dataset in RECV and compared generated rationales against those provided by human annotators. We restricted sample selection to those where at least three models predicted wrong veracity labels. We follow Song et al. (2024) and compute Factual Consistency (FC), Evidence Appropriateness (EA), BARTScore Yuan et al. (2021), and Perplexity (PPL) to assess the quality of the generated explanations. We provide additional details about metrics in Appendix G.
6.1 Results
VitaminC CLIMATE-FEVER PHEMEPlus Model EA $\uparrow$ FC $\uparrow$ BART $\uparrow$ PPL $\downarrow$ EA $\uparrow$ FC $\uparrow$ BART $\uparrow$ PPL $\downarrow$ EA $\uparrow$ FC $\uparrow$ BART $\uparrow$ PPL $\downarrow$ Claude ZS 0.85 0.85 -4.16 99.63 0.87 0.88 -4.26 31.08 0.82 0.81 -4.31 39.82 Claude ZS CoT 0.82 0.83 -4.38 52.85 0.82 0.83 -4.28 29.57 0.85 0.84 -4.44 38.81 Claude M-CoT 0.85 0.86 -4.05 68.53 0.89 0.90 -3.42 25.52 0.89 0.88 -4.17 37.83 GPT-4 ZS 0.87 0.87 -3.83 66.84 0.91 0.91 -3.67 27.93 0.87 0.86 -3.89 40.83 GPT-4 ZS CoT 0.87 0.86 -3.78 59.01 0.90 0.91 -3.65 20.65 0.87 0.87 -3.89 28.83 GPT-4 M-CoT 0.89 0.88 -2.98 45.84 0.93 0.94 -2.90 28.13 0.85 0.85 -3.40 47.15 GPT-4o ZS 0.90 0.88 -3.63 52.96 0.92 0.93 -3.64 58.42 0.85 0.86 -4.01 99.63 GPT-4o ZS CoT 0.91 0.89 -3.45 35.82 0.93 0.94 -3.39 46.92 0.89 0.90 -3.74 52.85 GPT-4o M-CoT 0.90 0.88 -3.63 50.10 0.90 0.91 -3.57 57.65 0.87 0.86 -4.08 68.53
Table 3: Qualitative evaluation on RECV in the Exp setting. Best results are in bold, second-best results are underlined.
Table 3 reports the results concerning explanation evaluation. We observe that all models achieve comparable results on appropriateness (EA), consistency (FC), and coherence measured via BARTScore (BART), while showing notable discrepancies regarding perplexity (PPL). In particular, GPT-4o ZS CoT has the most faithful rationales across all datasets. Moreover, prompting strategies like ZS CoT and M-CoT do not lead to consistent improvements over ZS, suggesting that their effectiveness may be problem- and model-dependent.
Additionally, we assess generated rationales regarding correct and wrong model predictions in Appendix G. Our results show that rationales from correct predictions better align with ground-truth explanations, suggesting that wrong predictions are usually the by-product of incorrect reasoning (the model is unable to leverage the explanation).
Lastly, we analyze how similar the generated rationales were between the models. To do so, we perform a permutation test using sentence-level contradiction scores from Fact_Score (see Appendix G). We find that Claude ZS has the most unique rationales on all datasets. We discuss properties of rationales generated in the case of abductive and deductive errors per dataset.
VitaminC
We observe that LLMs struggle to generate faithful rationales in abductive cases. In particular, models tend to generate assertions rather than hedged information. This has implications for claim verification, where models predominantly refute or misclassify the veracity of the claim based on the generated explanations. Regarding deductive reasoning, we observe that the majority of errors are due to internal biases of LLMs, heavily influencing rationale generation, and to semantic faults in attending to only some parts of the claim and evidence.
CLIMATE-FEVER
Regarding abductive cases, we observe the same issue reported in VitaminC. Regarding deductive reasoning, the majority of failures are due to implicit reasoning where relevant evidence information is implicit or where temporal relations between factual content must be understood to reach the correct conclusion.
PHEMEPlus
Contrary to VitaminC and CLIMATE-FEVER, abductive and deductive reasoning errors are mainly due to semantic interpretation issues where models focus only on specific information in the claim and evidence. This limits models in assessing claim-evidence pairs in their entirety, thus, hindering them in capturing relations between the evidence and the claim. As in VitaminC, this issue affects the claim verification performance, often leading to misclassification.
7 Findings
We discuss the main findings of our work, including task complexity, the effectiveness of prompting strategies, and rationale generation.
Reasoning and Task complexity.
Our results show that abductive reasoning is consistently more challenging than deductive reasoning. In particular, the performance gap between the two cases is around three times on average. This is mainly motivated by LLMs failing in performing uncertainty reasoning, often leading to erroneous assertive conclusions. Nonetheless, this is not the only issue that makes RECV challenging; task complexity plays a crucial role in reasoning performance. For instance, PHEMEPlus represents a more complex setting than VitaminC where news articles can contain extensive amount of information compared to Wikipedia pages. As shown in our qualitative analysis, LLMs tend to focus only on specific parts of input claim-evidence pairs, leading to suboptimal performance. For example, in the following claim from VitaminC,
Claim: Peking University is in a unitary sovereign state that’s located in East Asia.
Evidence: Peking University abbreviated PKU is a major Chinese research university located in Beijing and a member of the C9 League.
the veracity of is deductively refuted. However, all the models labelled this pair as evidence supporting the claim. The rationale provided was that China was a sovereign country, ignoring the claim completely. This shows the over-reliance on specific parts of the evidence by the models while ignoring others. Overall, our findings suggest that LLMs’ reasoning capabilities are domain and task dependent. Thus, we believe RECV represents a valuable resource to assess reasoning capabilities since it covers a wide spectrum of settings concerning claim verification.
Prompting Settings and Strategies
Our experiments show that prompting strategies like ZS CoT and M-CoT do not lead to systematic performance improvements, but are rather specific to datasets (e.g., VitaminC) and models. These results align with recent findings about CoT being beneficial mainly for math- and code-related tasks (Sprague et al., 2024). This is likely derived by divergent reasoning paths within the models during inference that lead to reduction in performance (Chollet, 2023; Todd et al., 2023; Dutta et al., 2024).
Furthermore, we also observe that internal alignment can hinder reasoning capabilities when it comes to abductive reasoning. Models are averse to provide predictions when evidence is incomplete. Yet abductive reasoning is often required for more complex tasks such as legal reasoning, just in time fact-checking, and other diverse forms of composite reasoning tasks. Hence, in order to achieve good results on these type of reasoning tasks, LLMs need to improve in the direction of abductive reasoning.
Explanation Quality
Our evaluation of generated explanations shows that these are on average consistent with human rationales. In particular, ZS CoT rationales are more convincing due to their verbosity, whereas M-CoT rationales are more concise. Moreover, we observe that rationales generated for abductive reasoning cases resemble assertions as models disprefer generating uncertain rationales. Nonetheless, considering that our results are limited to macro performance results and given the limited number of abductive cases, we believe our estimates to decrease as dataset size increases. We leave a fine-grained analysis on generated rationales concerning an extended version of RECV as future work.
8 Related work
LLMs for Reasoning
Several contributions have evaluated different reasoning capabilities in LLMs, including atomic and compounds types.For instance, LLMs can perform abductive reasoning for event prediction Shi et al. (2023), but struggle with common sense reasoning Zhao et al. (2023). Similarly, deductive reasoning in LLMs is beneficial to theorem-proving Saparov et al. (2023), factual content generation Akyürek et al. (2024), and question-answering Li et al. (2024). Nonetheless, the observed improvements are often attributable to how prompts are designed rather than an emergent deductive capability Chen et al. (2024). Moreover, LLMs perform out-of-context inductive reasoning Treutlein et al. (2024), but fail in lexical tasks Ye et al. (2023). Regarding analogical reasoning, LLMs address a wide variety of tasks, including nonverbal tests Webb et al. (2023); Hu et al. (2023), question-answering Yu et al. (2023), mathematical problem solving Yasunaga et al. (2024), and planning Yu et al. (2024), but present shortcomings in as many others Ye et al. (2024); Sourati et al. (2024); Stevenson et al. (2024); Ahrabian et al. (2024); Lewis and Mitchell (2024).
Likewise, despite promising results in compound reasoning tasks, such as counterfactual Wu et al. (2023), and compositional reasoning Lu et al. (2023a), LLMs are notably unreliable Gao et al. (2023); Zhang et al. (2024), sensitive to context Hosseini et al. (2024); Chang and Bergen (2024), and rely on shortcuts Yang et al. (2024).
LLMs for Claim Verification
Early work with LLMs focused on verifying simple facts Lee et al. (2020). More recently, LLMs for claim verification have been augmented with external knowledge Li et al. (2023a); Cheung and Lam (2023), prompt-based reasoning Cao (2023); Li et al. (2023b); Lin et al. (2023), claim decomposition for fine-grained search into text chunks Li et al. (2023a) or first-order logic terms Wang and Shu (2023), and data-augmentation Alhindi et al. (2023). While LLMs have been extensively applied in fact-checking, the question of which reasoning capabilities are needed to verify claims remains unaddressed. Thus, we are the first to propose a reasoning benchmark for claim verification.
9 Conclusion
We propose a novel extendable logical reasoning framework for deconstructing claim-evidence pairs into reasoning steps, required to determine the veracity of a claim. We use our framework to create RECV, the first reasoning benchmark for claim verification focussed on deductive and abductive reasoning. Our results show that LLMs notably struggle with abductive reasoning, while performing better in deductive cases. Our findings show that LLMs reasoning capabilities are domain and task dependent. In particular, no specific prompting strategy, including rationale generation, is systematically beneficial across all datasets and models. Nevertheless, rationales generated by LLMs for deductive reasoning are on average consistent with human ones. Overall, these results highlight that RECV represents a challenging reasoning setting for LLMs and further research is required to reach satisfying performance.
Limitations
Dataset Selection and Reasoning Types
Our focus on deductive and abductive reasoning types is dictated by our findings in the preliminary study. Nonetheless, other resources could be investigated to expand our approach to include other reasoning types. An example domain is biomedicine where datasets like COVID-Fact Saakyan et al. (2021) could include examples where inductive reasoning is required to infer claim veracity.
Models
We analyse three widely adopted proprietary LLMs. However, other models, including open-source ones, are also widely assessed in reasoning tasks. For a broader evaluation of LLMs, our study could include other models although these are currently unlikely to outperform the most established proprietary models.
Rationale Generation
When LLMs generate an explanation, there is no guarantee that it is true to the final label assigned by the model. We mitigate this issue by obtaining both the label and explanation in the same prompt, although it should still be treated as merely "a plausible post-hoc explanation generated by the model" rather than the specific reason behind its decision.
Ethics Statement
The PHEMEPlus dataset is a pre-existing dataset of rumours, for which ethical approval was obtained by the original research team. The rest of the datasets were sampled from pre-existing datasets for which no ethical approval was required.
Appendix A Logical Reasoning Examples
We hereby provide a formal description of atomic reasoning types, including examples on claim verification whenever possible.
Deductive.
Deductive reasoning or top-down logic is a logical reasoning process where we use inference rules such as modus ponens to deduce the veracity of a conclusion based on multiple hypotheses.
A core element of deductive inference is that if the premises are true, then the conclusion is true. In formal logic, the rules of deduction are infinite Morishita et al. (2023), where the most common ones are modus ponens, syllogism, and elimination. The reader is referred to the works of Morishita et al. (2023) and Saparov et al. (2023) for a more in-depth discussion of deduction rules.
Example.
  • Claim: Schools closed, Dammartin-en-Goele residents told to stay indoors, town ‘like warzone’.
  • Evidence: Schools went into lockdown and the town appealed to residents to stay inside resident’s houses.
  • Conclusion: The evidence references the school closing down and residents being told to shelter at home. Therefore, we deductively infer that the rumour is true as the conclusion logically follows the evidence.
Abductive.
There is much debate regarding definition of abductive reasoning (Plutynski, 2011). We follow the work of Paul (1993), which provides three different approaches towards defining abductive reasoning as:
  • A set-cover-based approach;
  • A logic-based approach;
  • A knowledge-level approach.
In this work, we use the set-cover-based approach, in which we construct the set of most plausible hypotheses $H$ given some observations $O$. Afterwards, we find the best possible explanation $E$ based on $H$. In other words,
  • A domain for hypothesis assembly is defined by the triple $\phi$, $\sigma$, $\epsilon$), where $\phi$ is a finite set of hypotheses, $\sigma$ is a set of observations and $\epsilon$ is a mapping from subsets of $\phi$ to subsets of $\sigma$. $\epsilon(\phi)$ is called the explanatory power of the set of hypotheses $\phi$ and determines the set of observations $\sigma$ accounts for. An assembly problem is given by a set $\sigma^{\prime}\subseteq\sigma$ of observations that have to be explained.(Paul, 1993).
Additionally, the key difference between abductive reasoning and the other forms of reasoning types is that, unlike the other types, abdctive reasoning works "backwards" towards the most plausible hypothesis from a given set of rules and happenings. Deductive reasoning is formulation of results based on rule and observation and inductive reasoning is formulation of rule based on result and observation. Whereas, abductive reasoning is formulation of an observation based on rule and result. For example from Flach and Kakas (2000):
Rule All the beans from this bag are white.
Result These beans are white.
Conclusion These beans are from this bag.
Inductive.
Inductive reasoning is the reasoning process where we use observations and outcomes to infer a generalizable rule. Hence, the logical structure can be represented as:
  • $\forall x,observations(x)\implies conclusion$
    or
  • $\exists x,observations(x)\implies conclusion$
amongst many other forms. A conclusion reached by inductive reasoning is not necessarily true. As per Flach and Kakas (2000), if the premises for any stated argument only provide partial support for its conclusion, then that argument is inductive supposing the premises are true.
Example 1.
  • Claim: Injecting or consuming bleach is good for killing the virus (Covid-19).
  • Evidence 1: Applying bleach or chlorine to the skin can cause harm, especially if it enters the eyes or mouth.
  • Evidence 2: These chemicals can disinfect surfaces, but people should not use them on their bodies.
  • Evidence 3: Also, these products cannot kill viruses inside the body.
  • Conclusion: From the evidence we can inductively draw a general conclusion that the claim is false, as bleach causes harm to the body and would not kill any viruses within.
Example 2.
Observation1 Eagles have wings. Eagles are birds and eagles can fly.
Observation2 Ducks have wings. Ducks are birds and ducks can fly.
Observation 3a Pigeons have wings. Pigeons are birds and pigeons can fly.
or
Observation 3b Bats have wings. Bats are mammals and bats can fly.
Conclusion a All birds have wings and all birds can fly.
or
Conclusion b Those who have wings can fly.
It is clear that each conclusion is true if we make a closed-world assumption regarding the premises. However, in reality, it is false as there exist flightless birds including Penguins and wingless birds such as Kiwi.
Analogical.
Analogical reasoning is the reasoning process concerned with comparison between two or more objects, arguments, or entities.
Example.
  • Claim: entity $\alpha$ is equivalent to entities $\zeta$, $\kappa$, $\phi$, and $\omega$.
  • Evidence: entity $\beta$ is equivalent to entities $\zeta$, $\kappa$, and $\phi$.
  • Conclusion: entity $\beta$ is probably equivalent to entity $\omega$.
Appendix B Claim Verification Datasets
We select three popular resources for claim verification, covering different domains and increasing task complexity.
VitaminC Schuster et al. (2021).
A multi-task fact-checking dataset based on manual and synthetic English revisions to Wikipedia pages. The dataset comprises $\sim$450k claim-evidence pairs. For the claim verification task, claim-evidence pairs are annotated with veracity labels: supports, refutes, and not-enough-information. VitaminC is licensed under MIT License.
CLIMATE-FEVER Diggelmann et al. (2020) .
A claim verification dataset that consists of $\sim$1.5k real-world claims concerning climate change. The claims are retrieved from Google while the evidence is Wikipedia-based. The claim-evidence pairs are annotated with veracity labels: supports, refutes, and not-enough-information.
PHEMEPlus Dougrez-Lewis et al. (2022).
A rumour verification dataset comprising social media claims about real-world events. The dataset contains five different events where associated claim-evidence pairs are annotated with veracity labels: true, false, not-enough-information. PHEMEPlus is an extension of the PHEME Zubiaga et al. (2016) dataset, where web-retrieved news articles are used as evidence in place of Twitter threads as done in PHEME.
Appendix C Sampling details
[…]
Figure A1: Distribution of the sampling metrics.
Appendix D Annotation Guidelines
Figure A2 summarizes our annotation pipeline for RECV. In data annotation (Figure A2, Top), we provide a human annotator with claim-evidence pairs with corresponding veracity label. The annotator determines the reasoning type required to infer the claim veracity and provides a rationale in free-text format as motivation.
Figure A2: (Top) Our annotation process for reasoning-based claim verification. An annotator provides reasoning type required to infer the claim veracity and a rationale as motivation. (Bottom) The claim verification task where a LLM has to predict the claim veracity and generate a rationale as support.
Table A1 reports the annotation guidelines we used to instruct annotators in creating RECV.
Read each claim and evidence pair samples along with their associated veracity labels. Afterwards you will label them with the type of reasoning you think was necessary for inferring the veracity label of the claim given the evidence. The reasoning types are abductive and deductive. Also, provide rationale for your labels.
The goal is to identify what type of reasoning is necessary to infer the veracity of the claim given the associated evidence, for each of the given pairs.
Example 1.
Claim: Climate change isn’t increasing extreme weather damage costs.
Evidence: 1. Many analyses, such as that of the Stern Review presented to the British Government, have predicted reductions by several percent of world gross domestic product due to climate related costs such as dealing with increased extreme weather events and stresses to low-lying areas due to sea level rises. 2. Global losses reveal rapidly rising costs due to extreme weather-related events since the 1970s. 3. Global warming boosts the probability of extreme weather events, like heat waves, far more than it boosts more moderate events. 4. "Impacts [of climate change] will very likely increase due to increased frequencies and intensities of some extreme weather events".
Veracity: Refutes
Reasoning: Deductive
Rationale: The evidence deductively refutes the claim. We find explicit mention of increased damage cost in the second line of the evidence. While the last two lines of evidence provide explicit evidence of global causing more adverse weather events.
Example 2.
Claim: Pluto’s climate change over the last 14 years is likely a seasonal event.
Evidence: The long orbital period of Neptune results in seasons lasting forty years. 2. As a result, Neptune experiences similar seasonal changes to Earth. 3. "Evidence for methane escape and strong seasonal and dynamical perturbations of Neptune’s atmospheric temperatures". 4. Each planet therefore has seasons, changes to the climate over the course of its year.
Veracity: Supports
Reasoning: Abductive
Rationale: The claim is abductively supported. Given Pluto used to be a planet and now is labeled as a dwarf planet, we can hypothesize that it likely has the same attribute as neptune. Given pluto has the biggest orbital period, it is very much likely pluto seasons last over 10 years.
Table A1: Annotation guidelines used to create RECV. This specific example was for CLIMATE-FEVER. This partial representation of the guideline as we provided 8 examples for each dataset with a mix of supports and refutes label.
Appendix E Data Annotation
Table A2 reports pairwise agreement scores for each dataset in RECV.
VitaminC CLIMATE-FEVER PHEMEPlus Annotator A Annotator B Annotator C Annotator A Annotator B Annotator C Annotator A Annotator B Annotator C Annotator A - 0.72 0.74 - 0.56 0.56 - 0.64 0.68 Annotator B 0.72 - 0.78 0.56 - 0.56 0.64 - 0.68
Table A2: Pairwise Bennett’s S Score across different datasets.
Appendix F Prompts
Tables A3, A4, and A5 report the prompts we used for VitaminC, CLIMATE-FEVER, and PHEMEPlus, respectively. We follow standard prompt construction strategies and provide dataset specific personas and instructions. Additionally, in M-CoT with provide examples to guide the model.
ZS.
[…]
CoT.
[…]
M-CoT.
[…]
Here are some examples:
{examples}
Table A3: Prompts used in VitaminC.
ZS.
[…]
CoT.
[…]
M-CoT.
[…]
Here are some examples:
{examples}
Table A4: Prompts used in CLIMATE-FEVER.
ZS.
[…]
CoT.
[…]
M-CoT.
[…]
Here are some examples:
{examples}
Table A5: Prompts used in PHEMEPlus.
Appendix G Qualitative Analysis
Tables A9, A10, and A11 report pairwise permutation tests on RECV datasets. Moreover, Tables A6, A7, and A8 report qualitative analysis metrics on RECV datasets. In particular, we compute qualitative metrics on two sets of examples: those for which models correctly predicted the corresponding claim veracity (Correct), and those where models made wrong predictions (Wrong).
The metrics used for qualitative analysis are as following,
Factual consistency.
We assess the consistency of LLM generated rationales $R$ with human-written ones $H$, where consistency is the absence of contradiction. We define $C$ to be a function that quantifies the consistency of text $B$ based on text $A$:
$C(A,B)=\frac{1}{|A|\cdot|B|}\sum_{a\in A}\sum_{b\in B}\left(1-\text{NLI}(\text{Contradict}|a,b)\right)$
We calculate the consistency of LLM rationales to human rationales as,
[…]
where $N$ is the total number of sentence pairs compared,$C_{i}$ is the consistency score of the $i$-th comparison.
Evidence appropriateness.
For evidence appropriateness, we use the same consistency score $C$ as Fact_Expert.
$\text{EA}=\frac{1}{M}\sum_{j=1}^{M}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}(1-c_{ij})\right)$ (2)
Here, $M$ is the total number of generated rationales, $N_{j}$ is the number of sentences in the $j$-th generated rationale and $C_{i}j$ is the consistency score for the $i$-th sentence in the $j$-th rationale. Evidence appropriateness can be considered as the mean factual consistency whereas Fact_Expert is the granular sentence level consistency.
Coherence.
We estimate how easy it is to follow the rationales and how effectively it integrates information from the evidence using BARTScore.
Fluency.
We estimate fluency for rationales using perplexity (PPL) under GPT-2-xl Radford et al. (2019).
EA $\uparrow$ FC $\uparrow$ BART $\uparrow$ PPL $\downarrow$ Model Correct Wrong Correct Wrong Correct Wrong Correct Wrong Claude ZS 0.86 0.75 0.86 0.77 -4.15 -4.21 103.49 75.94 Claude ZS CoT 0.86 0.61 0.86 0.75 -4.42 -4.19 55.44 39.25 Claude M-CoT 0.87 0.82 0.86 0.84 -4.08 -3.90 72.02 47.03 GPT-4 ZS 0.89 0.76 0.88 0.82 -3.82 -3.87 69.82 51.17 GPT-4 ZS CoT 0.89 0.76 0.88 0.80 -3.77 -3.8 61.13 47.85 GPT-4 M-CoT 0.90 0.77 0.90 0.75 -2.98 -3.01 46.39 40.95 GPT-4o ZS 0.91 0.79 0.90 0.81 -3.67 -3.41 57.60 24.46 GPT-4o ZS CoT 0.93 0.75 0.92 0.76 -3.47 -3.29 37.77 21.57 GPT-4o M-CoT 0.92 0.77 0.90 0.78 -3.68 -3.35 53.63 26.52
Table A6: Qualitative evaluation on VitaminC. We distinguish between correct and wrong claim veracity predictions.
EA $\uparrow$ FC $\uparrow$ BART $\uparrow$ PPL $\downarrow$ Model Correct Wrong Correct Wrong Correct Wrong Correct Wrong Claude ZS 0.87 0.88 0.87 0.89 -4.19 -4.60 61.26 79.73 Claude ZS CoT 0.81 0.86 0.91 0.86 -4.29 -4.26 29.27 25.08 Claude M-CoT 0.90 0.82 0.91 0.85 -3.36 -3.68 33.51 33.33 GPT-4 ZS 0.93 0.79 0.93 0.84 -3.65 -3.79 30.17 36.24 GPT-4 ZS CoT 0.92 0.81 0.92 0.87 -3.63 -3.76 29.23 31.51 GPT-4 M-CoT 0.96 0.70 0.96 0.75 -2.88 -3.07 25.50 25.65 GPT-4o ZS 0.93 0.89 0.93 0.90 -3.63 -3.71 28.09 26.62 GPT-4o ZS CoT 0.95 0.85 0.95 0.90 -3.37 -3.48 20.77 19.95 GPT-4o M-CoT 0.90 0.88 0.91 0.86 -3.56 -3.72 28.15 28.04
Table A7: Qualitative evaluation on CLIMATE-FEVER. We distinguish between correct and wrong claim veracity predictions.
EA $\uparrow$ FC $\uparrow$ BART $\uparrow$ PPL $\downarrow$ Model Correct Wrong Correct Wrong Correct Wrong Correct Wrong Claude ZS 0.85 0.76 0.84 0.74 -4.29 -4.38 60.15 53.23 Claude ZS CoT 0.86 0.81 0.85 0.81 -4.43 -4.49 48.25 42.95 Claude M-CoT 0.89 0.87 0.89 0.86 -4.11 -4.30 57.30 58.42 GPT-4 ZS 0.88 0.81 0.89 0.81 -3.90 -3.87 41.33 35.52 GPT-4 ZS CoT 0.88 0.85 0.88 0.85 -3.90 -3.85 39.65 36.18 GPT-4 M-CoT 0.91 0.70 0.89 0.72 -3.40 -3.39 40.59 30.73 GPT-4o ZS 0.90 0.74 0.90 0.75 -4.02 -4.00 43.29 34.20 GPT-4o ZS CoT 0.92 0.82 0.92 0.84 -3.76 -3.70 30.33 24.55 GPT-4o M-CoT 0.89 0.79 0.89 0.79 -4.08 -4.05 49.46 40.59
Table A8: Qualitative evaluation on PHEMEPlus. We distinguish between correct and wrong claim veracity predictions.
Claude ZS Claude ZS CoT Claude M-CoT GPT-4 ZS GPT-4 ZS CoT GPT-4 M-CoT GPT-4o ZS GPT-4o ZS CoT Claude ZS CoT 0.4061 - - - - - - - Claude M-CoT 0.5985 0.1768 - - - - - - GPT-4 ZS 0.3461 0.0639 0.6466 - - - - - GPT-4 ZS CoT 0.3303 0.0600 0.6636 0.9703 - - - - GPT-4 M-CoT 0.0830 0.0066 0.1943 0.3769 0.3489 - - - GPT-4o ZS 0.0732 0.0067 0.1897 0.4086 0.3739 0.8896 - - GPT-4o ZS CoT 0.0158 0.0002 0.0558 0.1511 0.1269 0.6856 0.5457 - GPT-4o M-CoT 0.0866 0.0064 0.2211 0.4259 0.3837 0.8893 0.9988 0.5626
Table A9: Pairwise Permutation Test on 100 evaluation samples from VitaminC.
Claude ZS Claude ZS CoT Claude M-CoT GPT-4 ZS GPT-4 ZS CoT GPT-4 M-CoT GPT-4o ZS GPT-4o ZS CoT Claude ZS CoT 0.0480 - - - - - - - Claude M-CoT 0.4062 0.0027 - - - - - - GPT-4 ZS 0.0811 0.0001 0.5795 - - - - - GPT-4 ZS CoT 0.0889 0.0002 0.6516 0.8770 - - - - GPT-4 M-CoT 0.0022 0.0001 0.0802 0.1529 0.1005 - - - GPT-4o ZS 0.0108 0.0001 0.2442 0.4047 0.3009 0.5190 - - GPT-4o ZS CoT 0.0006 0.0001 0.0528 0.0932 0.0536 0.9319 0.4036 - GPT-4o M-CoT 0.1325 0.0001 0.7312 0.7618 0.8751 0.0741 0.2653 0.0369
Table A10: Pairwise Permutation Test on 100 evaluation samples from CLIMATE-FEVER.
Claude ZS Claude ZS CoT Claude M-CoT GPT-4 ZS GPT-4 ZS CoT GPT-4 M-CoT GPT-4o ZS GPT-4o ZS CoT Claude ZS CoT 0.1872 - - - - - - - Claude M-CoT 0.0068 0.0953 - - - - - - GPT-4 ZS 0.0276 0.3199 0.4636 - - - - - GPT-4 ZS CoT 0.0082 0.1534 0.6888 0.7094 - - - - GPT-4 M-CoT 0.2363 0.8827 0.2256 0.4959 0.3116 - - - GPT-4o ZS 0.0534 0.4611 0.3528 0.8139 0.5373 0.6487 - - GPT-4o ZS CoT 0.0002 0.0033 0.4062 0.0791 0.1584 0.0263 0.0518 - GPT-4o M-CoT 0.0270 0.3188 0.4208 0.9707 0.6609 0.5043 0.8387 0.0681
Table A11: Pairwise Permutation Test on 100 evaluation samples from PHEMEPlus.